OpenStudy (anonymous):

the measures of the lengths of two sides of an isosceles triangle are x and 2x+1. Then, the perimeter of the triangle is (A) 4x (B) 4x+1 (C) 5x+1 (D) 5x+2 (E) None of the above

OpenStudy (anonymous):

An isoceles triangle has atleast two sides equal. So the perimeter is either: x+x+2x+1 = 4x+1 or x+2x+1+2x+1 = 5x+2 So the answer is B and D

OpenStudy (anonymous):

It would actually need to be just D; by the triangle inequality (the lengths of any two sides must be greater than the length of the third side) the longer, and therefore 'double' side would be of length 2x + 1. Thus, the answer would be 2 (2x + 1) + x = 5x + 2.

myininaya (myininaya):

2x+1>x for any x>0 so wouldn't it be x+x+2x+1=4x+1